A machine has an effiency of 75% the work done when the machine is used to raise a load of 50kg through a vertical distance of 6m is

1
Expert's answer
2022-01-28T12:50:52-0500

Given:

"\eta=0.75"

"m=50\:\rm kg"

"h=6\:\rm m"


The work done by the machine

"W=\frac{mgh}{\eta}=\rm\frac{50*9.8*6}{0.75}=3920\:J"